Hi this is a question in my summer review packet. I have forgotton how to figure it out if you could help that would be great.
The long leg of a 30-60-90 triangle is 3 inches. find the measure of the other 2 sides.

Answer by checkley71(8403)

the sides of a 30,60,90 right triangle are x,2x & xsqrt3 the longest leg is 2x=3 in. & the shortest leg is x or 2x=3 thus x=3/2 or x=1.5 in. and the third leg is 1.5sqrt3 or 1.5*1.732=2.59 in.